R41070

Yarragadee 12m

Disk VSN: HOB+0015/6000/1024 | Data volume at beginning: …. GB

  • 1830 UT Experiment started OK (EARL)
  • 1821 UT Modules were swapped for new experiment aum055 before old experiment ended. So last two scans are on new disk HOB+0064 (David Schunck)